40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 30 15:22:17 2011, Katrin, Update, Green Locking, What happened on green YARM? 
    Reply  Fri Sep 30 16:17:10 2011, kiwamu, Update, Green Locking, things to be done for the Yarm green lock 
Message ID: 5586     Entry time: Fri Sep 30 16:17:10 2011     In reply to: 5585
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: things to be done for the Yarm green lock 

Thank you for the summary.

I think now you are getting into a phase where you should start doing some quantitative and careful checks.

    1. Calculate how much light will be reflected from the cavity if the alignment is perfect.

    2. Investigate where those kHz oscillations are coming from.

    3. Estimate how much the 1.1 kHz corner frequency in LPF will reduce the phase margin around 10 kHz.

    4. Calculate the estimated amplitude of the PDH signal.

    5. Compute how big the gain of the PDH box should be.

Quote from #5585

 This is a kind of summary of what I have worked on this week. 

  • DC level of green PD when light resonates 75%
                    --> Not sure if this alignment is good enough
  • The PDH error signal was covered with oscillations of 3.3 kHz, 7.1 kHz and 35 kHz.
  • Designed and build a new LPF: second order, cut of frequency of 1.1 kHz (this is just the design value, haven't measured that so far)
  • Signal-to-noise ratio (SNR) could be improved to values between 7.8 and 11.1 (old SNR was 5 to 7)

ELOG V3.1.3-